Nitrogen monoxide reacts with hydrogen to form nitrogen and water, as shown by the following balanced chemical equation.
2NO (g)+ 2H2(g) = N2 (g) + 2H2O (g)
A scientist plans to react 9.0 x 10-2 L NO held at a pressure of 200.0 kPa and a temperature of 20.0 °C with excess hydrogen. How many moles of nitrogen gas will be produced?
1
Expert's answer
2020-05-05T03:37:51-0400

Using the Clapeyron-Mendeleev equation: "PV=nRT"  , P - gas pressure (Pa) , V - gas volume (m3), n - chemical amount of gas (mol), T - temperature of gas (K), R - Gas constant (8,314 J/(mol*K)).

Express n: "n=PV\/RT"

P=200,0kPa=200000Pa

T=20,0oC=293,15K

V=9.0 x 10-2 L=9*10-5m3

n(NO)=200000Pa*9*10-5m3/(8,314 J/(mol*K)*293,15K)=0,0074mol

According to the reaction equation n(N2)=n(NO)/2=0,0074mol/2=0,0037mol

Answer:0,0037mol